An ideal spring hangs from the ceiling. A 1.45 kg mass is hung from the spring, stretching the spring a distance d = 0.0865 m from its original length when it reaches equilibrium. The mass is then lifted up a distance L = 0.0275 m from the equilibrium position and released. What is the kinetic energy of the mass at the instant it passes back through the equilibrium position?
Answer #1

At equilbrium

Elongation =d

So kd=mg

Angular frequency

\omega =\sqrt{\frac{k}{m}}

10.644rad/s

Amplitude of SHM =L=0.0275m

So

Speed at equilibrium position = wA=wL=10.644×0.0275=0.273m/s

Kinetic.energy K=0.5mv^2=0.062J